OpenStudy (anonymous):

Giving medal and fan if you solve! Tina described four triangles as shown below: Triangle A: All sides have length 12 cm. Triangle B: Two sides have length 10 cm, and the included angle measures 60°. Triangle C: Base has length 15 cm, and base angles measure 40°. Triangle D: All angles measure 60°. Which triangle is not a unique triangle?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (michele_laino):

I think D. because there is an infinity of triangles which have all three angles, whose amplitude is 60°, namely those triangles are similar each other

OpenStudy (anonymous):

thx need with another one can you help!! :)

OpenStudy (michele_laino):

I'm ready!

OpenStudy (anonymous):

What is the value of x? http://learn.flvs.net/webdav/assessment_images/educator_mjmath2_v14/05_14_p1_20.jpg

OpenStudy (michele_laino):

I can't access, may be because I'm not a student!

OpenStudy (anonymous):

I can draw it

OpenStudy (michele_laino):

Ok!

OpenStudy (anonymous):



OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

from your drawing, the sum of the amplitudes of the two angles, namely 2x+10, must be equals to 90°, because, that sum between the two angles is equal to a right angle, so: \[2x+10=90\] ok!, now adding -10 to both sides of last equation, we have: \[2x+10-10=90-10\] or: \[2x=80\] finally if I divide the above last equation by 2, I get: \[\frac{ 2x }{ 2 }=\frac{ 80 }{ 2 }\] so: \[x=40°\]

OpenStudy (anonymous):

thx so much! You have a new fan

OpenStudy (michele_laino):

thank you!

OpenStudy (anonymous):

A running track in the shape of an oval is shown. The ends of the track form semicircles. A running track is shown. The left and right edges of the track are identical curves. The top and bottom edges of the track are straight lines. The track has width 56 m and length of one straight edge 130 m. What is the perimeter of the inside of the track? (π = 3.14)

OpenStudy (anonymous):

@Michele_Laino